### v1

@@ Abstract
Potato yield is an important metric for farmers to further optimize their cultivation practices. Potato yield can be estimated on a harvester using an RGB-D camera that can estimate the three-dimensional (3D) volume of individual potato tubers. A challenge, however, is that the 3D shape derived from RGB-D images is only partially completed, underestimating the actual volume. To address this issue, we developed a 3D shape completion network, called CoRe++, which can complete the 3D shape from RGB-D images. CoRe++ is a deep learning network that consists of a convolutional encoder and a decoder. The encoder compresses RGB-D images into latent vectors that are used by the decoder to complete the 3D shape using the deep signed distance field network (DeepSDF). To evaluate our CoRe++ network, we collected partial and complete 3D point clouds of 339 potato tubers on an operational harvester in Japan. On the 1425 RGB-D images in the test set (representing 51 unique potato tubers), our network achieved a completion accuracy of 2.8 mm on average. For volumetric estimation, the root mean squared error (RMSE) was 22.6 ml, and this was better than the RMSE of the linear regression (31.1 ml) and the base model (36.9 ml). We found that the RMSE can be further reduced to 18.2 ml when performing the 3D shape completion in the center of the RGB-D image. With an average 3D shape completion time of 10 milliseconds per tuber, we can conclude that CoRe++ is both fast and accurate enough to be implemented on an operational harvester for high-throughput potato yield estimation. Our code, network weights and dataset are publicly available at https://github.com/UTokyo-FieldPhenomics-Lab/corepp.git.
@@ 1 Introduction
Potatoes (Solanum tuberosum) are important for global food security and can complement a cereal-based diet due to their high nutritional value and high yield (Zhang et al., 2017). To further optimize potato yields, farmers must have sub-field level information on field yields at a 5-10 meter scale, as this provides important insights into the marketable yield of the field, as well as to further optimize the fertilization and cultivation practices. Sub-field level potato yield monitoring can be performed by systems installed on the harvester that can automatically estimate key yield parameters such as tuber number, size, volume, and weight. A commonly used potato yield monitoring system uses load cells attached to the harvesters’ conveyor belt to measure the mass of the harvested produce in real time (Zamani et al., 2014; Kabir et al., 2018). Such weighing systems are simple to use and maintain, but a disadvantage of these systems is that tare (such as stones, soil and crop residue) is also included in the weight measurement, making the tuber yield profile unobtainable.
@@
A more promising alternative is to use a camera system that can visually distinguish potato tubers from tare, allowing the system to count and measure only the potato tubers. Currently, at least 14 of such camera systems have been presented in scientific literature: Noordam et al. (2000); Hofstee and Molema (2003); ElMasry et al. (2012); Razmjooy et al. (2012); Lee et al. (2018); Long et al. (2018); Si et al. (2018); Su et al. (2018); Pandey et al. (2019); Cai et al. (2020); Lee and Shin (2020); Dolata et al. (2021); Huynh et al. (2022); Jang et al. (2023). Most of these systems (11 of the 14) used a single red, green and blue (RGB) color camera to detect the potato tubers and then estimate their yield parameters using offline calibrated pixel-to-world conversion factors. While this method proved effective in calibrated laboratory setups, it was also found to have limited accuracy for potatoes that are occluded. Occlusion is a common phenomena on the conveyor belt of an operational harvester.
@@
More recent studies have focused on extending the traditional RGB camera with additional three-dimensional (3D) vision. With 3D vision, it is possible to create partial or complete 3D shapes of potato tubers, allowing for better estimation of tuber yield under the challenging conditions of occlusion. Cai et al. (2020) extended an RGB camera with a laser line triangulation method to capture the complete 3D shape of potato tubers. Although their system enabled very accurate volumetric estimates of potato tubers in a laboratory setup, the image acquisition method was unfortunately too time-consuming to be applied on an operational harvester. RGB cameras with embedded depth sensing abilities, known as RGB-D cameras, allow much higher throughput than laser triangulation methods. RGB-D cameras, such as the ones used by Long et al. (2018) and Su et al. (2018), are therefore more promising for use on an operational harvester.
@@
A current limitation with RGB-D cameras is that they can only produce a partial 3D shape of the potato tuber, which can lead to an underestimation of the actual size, volume or weight. Combining multiple RGB-D cameras can potentially reduce the effect of this problem, but this will not fully solve it and make the overall system complex and more expensive. Therefore, it is more desirable and economically feasible to use a single RGB-D camera, and then estimate the complete 3D shape with dedicated software. There are numerous examples in the scientific literature of using 3D shape completion software to estimate complete shapes from partial shapes. Most of the current shape completion methods use deep learning techniques based on multi-layered perceptrons, graph-based convolutional neural networks and encoder-decoder networks to complete the 3D shape from partially completed point clouds (Fei et al., 2022). Also in the agricultural domain, there are a few studies on 3D shape completion: one on the completion of plant leaves (Chen et al., 2023), one on the completion of trees (Xu et al., 2023), and five on the completion of fruits (Ge et al., 2020; Magistri et al., 2022, 2024; Marangoz et al., 2022; Pan et al., 2023).
@@
When assessing the applicability of 3D shape completion methods for potato yield estimation, it is crucial to consider their processing speed. Ideally, the shape completion method should be fast enough to process all the potato tubers that move over the conveyor belt during harvest, because this will provide the most complete and accurate yield estimate possible. Since roughly more than 200,000 potato tubers are harvested per hectare, the 3D shape completion method must complete processing in a matter of milliseconds. Of the studies mentioned above, only the method of Magistri et al. (2022) has the potential to do so. Magistri’s method, called CoRe (Completion and Reconstruction), can complete 3D shapes in 4 milliseconds (tested on sweet peppers and strawberries). The novelty of CoRe is the addition of a convolutional encoder to the deep signed distance function decoder (DeepSDF, Park et al., 2019). CoRe’s convolutional encoder compresses RGB-D images into a latent vector, which is a compact representation of the shape’s geometry. The latent vector from the encoder is then used along with the 3D coordinates of the point cloud as input for DeepSDF to reconstruct the complete 3D shape. Because the latent vector is predicted by the convolutional encoder, there is no need for DeepSDF’s original time-consuming latent vector optimization, making the entire 3D shape completion method significantly faster.
@@
The main contribution of this paper is the extension of the work by Magistri et al. (2022) to answer these research questions. We have conducted a systematic study and analysis on how to optimize this 3D shape completion method. Our optimization has led to several improvements to CoRe’s original convolution encoder: an improved data preprocessing, geometric and color data augmentations, an updated loss function, an updated neural network architecture, and a graphical processing unit (GPU)-based 3D mesh generation that improves and speeds up the 3D shape completion.
This paper presents the research and development of a 3D shape completion network for estimating the volume of individual potato tubers on an operational harvester. Our research novelties are four-fold. First, we optimized CoRe’s original convolutional encoder for faster and more accurate 3D shape completion of potato tubers on an operational harvester. Second, we performed a systematic analysis of the effect of the latent size, image analysis region, potato size, and potato cultivar on the performance of the 3D shape completion of potato tubers. Third, we conducted two ablation studies on the impact of our new additions to the overall performance. Fourth, we publicly released our 3D dataset with partially and fully completed point clouds of potato tubers collected on an operational harvester in dirty and cluttered circumstances.
@@ 2.1.1 Imaging system
We installed an imaging system above the conveyor belt of a single-row potato harvester (Toyonoki Top-1, Figure 1a). The imaging system was a black plastic box in which four light emitting diode (LED) strips were installed to provide light (Figure 1b). Reflective curtains on the sides of the plastic box helped to diffuse the light (Figure 1c). In the center of the box, an RGB-D camera (Intel Realsense D405) was installed. The distance between the RGB-D camera and the conveyor belt was approximately 0.33 m. At this distance, the camera’s field of view was 0.64 m (width) by 0.39 m (height). The RGB-D camera was connected to a laptop computer (Lenovo P51) that used the Robot Operating System 2 (ROS2, version: Humble Hawksbill) to collect color and depth images with 30 frames per second in ROS2 bag files.
Figure 1: (a) Overview of the imaging system installed on a potato harvester in Sarabetsu, Japan. (b) Close-up photo of the imaging system on the harvester. (c) Inside the imaging box, an RGB-D camera (Intel Realsense D405) was installed. Four LED strips provided the necessary illumination inside the box. The sides of box were covered with a reflective curtain that provided diffuse lighting conditions.
@@ 2.1.2 Image collection on the harvester
While running the harvester in the field or in the barn, we collected images of 339 potato tubers of different sizes and shapes to test our 3D shape completion method. Our collection method consisted of the following procedure: one person standing in front of the imaging box selected a potato tuber of a random size and shape and then inserted a colored thumbtack into the potato. The thumbtack was inserted into the potato such that it was visible in the RGB image when the tuber passed the image acquisition region of the RGB-D camera. Given the speed of the conveyor belt, we were able to capture between 20 and 30 images of the marked potato tuber as it moved under the camera along with the conveyor belt (Figure 2). After the image acquisition, another person standing behind the imaging box collected the marked potato by hand. The marked potato was then placed in a bucket that was later brought to the barn for 3D reconstruction.
@@
Figure 2: Our image collection method involved marking potato tubers with a colored thumbtack so that the tuber could be easily identified in the image and easily collected after image acquisition. (a) In this example, the tuber with the red thumbtack was first fully photographed when it entered the image acquisition area of the RGB-D camera in the bottom of the image. (b) This is the 15th captured image in which the marked potato is in the center of the image. (c) This is the 29th and last captured frame in which the marked tuber was fully photographed.
2.1.3 3D reconstruction of the collected potato tubers
@@
To obtain the complete 3D shape of the collected potato tubers, we have set up an imaging system in the barn (Figure 3a). Our imaging system consisted of three Canon X7 DSLR cameras, a turntable, four auto-detectable marker stands (Fig. 3b), and a photo studio with LED illumination.
@@
Prior to taking the images, each collected potato tuber was pierced with a narrow threaded bolt attached to a white wooden board (Figure 3b). This piercing allowed the tuber to be photographed from three camera perspectives while being held off the ground, allowing almost the entire longitudinal section of the tuber to be photographed at once. To capture the different sides of the tuber, the turntable was set to automatically rotate 15 degrees and then stop for two seconds to give the three cameras time to acquire the images. This was repeated until the tuber was photographed from all sides. The image acquisition was accomplished with a commercial shutter controller (Esper TriggerBox) that released an electronic trigger to the three cameras at the same time. The images were directly transferred to the connected laptop computer to store them for later 3D reconstruction (Figure 3c). The camera parameter settings, image transferring, and image renaming were accomplished with DigiCamControl software, of which its details are shown in Figure 3d.
Figure 3: The workflow of our 3D reconstruction included three steps: (1) image collection, (2) image preprocessing, (3) 3D reconstruction with Structure-from-Motion (SfM). (a) Our image acquisition hardware consisted of three Canon DSLR cameras attached to a frame, a turntable, four auto-detectable marker stands, and a photo studio with LED illumination. (b) Before image acquisition, a threaded bolt was inserted into the potato tuber such that it was lifted off the base. This allowed the three cameras to photograph an entire longitudinal section of the tuber at once. (c) A multi-camera control trigger box was used to acquire images from the three cameras at the same time. (d) DigiCamControl software was used to adjust the cameras’ parameters. (e-h) After image acquisition, the region of the potato tuber was extracted through CIELAB thresholding (f) and deep learning refinement (g). (i) Metashape batch processing was used to reconstruct the 3D shape of the potato tubers.
After the image acquisition, the images were processed to extract the region of the potato tuber. This region extraction was accomplished by applying a threshold on the CIELAB color space to filter out the white- and black-colored background (Figure 3e) (the threshold was set to larger than 15 for the A channel and B channel). The resulting mask was refined by using the pre-trained CascadePSP (Cheng et al., 2020) deep learning network (Figure 3g&h).
@@
After generating the masks of all tubers, we implemented an automatic batch processing workflow based on the Metashape API for 3D reconstruction (Figure 3i). First, the images and corresponding masks of each potato tuber were grouped for each camera and its corresponding view angles. Then, the control point markers in the images were automatically detected, the scale bars were automatically imported, and world coordinates were automatically assigned to each potato tuber. Afterwards, the images of the different cameras were aligned, and the corresponding key points were generated. These aligned images served as input for the 3D reconstruction, which resulted the 3D mesh model, including its colored model textures. After 3D reconstruction, small distortions, such as 3D meshes belonging to the white wooden board were removed. Then, the filtered meshes were double-checked by another researcher in CloudCompare who removed small disconnected components if necessary. Figure 4 illustrates a filtered 3D mesh with colored textures of a potato tuber.

### v3

@@ Abstract
Potato yield is an important metric for farmers to further optimize their cultivation practices. Potato yield can be estimated on a harvester using an RGB-D camera that can estimate the three-dimensional (3D) volume of individual potato tubers. A challenge, however, is that the 3D shape derived from RGB-D images is only partially completed, underestimating the actual volume. To address this issue, we developed a 3D shape completion network, called CoRe++, which can complete the 3D shape from RGB-D images. CoRe++ is a deep learning network that consists of a convolutional encoder and a decoder. The encoder compresses RGB-D images into latent vectors that are used by the decoder to complete the 3D shape using the deep signed distance field network (DeepSDF). To evaluate our CoRe++ network, we collected partial and complete 3D point clouds of 339 potato tubers on an operational harvester in Japan. On the 1425 RGB-D images in the test set (representing 51 unique potato tubers), our network achieved a completion accuracy of 2.8 mm on average. For volumetric estimation, the root mean squared error (RMSE) was 22.6 ml, and this was better than the RMSE of the linear regression (31.1 ml) and the base model (36.9 ml). We found that the RMSE can be further reduced to 18.2 ml when performing the 3D shape completion in the center of the RGB-D image. With an average 3D shape completion time of 10 milliseconds per tuber, we can conclude that CoRe++ is both fast and accurate enough to be implemented on an operational harvester for high-throughput potato yield estimation. CoRe++’s high-throughput and accurate processing allows it to be applied to other tuber, fruit and vegetable crops, thereby enabling versatile, accurate and real-time yield monitoring in precision agriculture. Our code, network weights and dataset are publicly available at https://github.com/UTokyo-FieldPhenomics-Lab/corepp.git.
@@ 1 Introduction
Potatoes (Solanum tuberosum) are important for global food security (Zhang et al., 2017) and are of considerable economic importance. In 2021, potato production contributed \$100 billion to the US economy (Knudson and Miller, 2023), and €19.4 billion to the EU economy in 2023 (Eurostat, 2024). Potatoes have relatively high nutrient requirements and the quantity and timing of nutrient application impacts tuber yield (Ruark et al., 2014; Love et al., 2005). Excess nutrient application represents wasted money and can contribute to waterway eutrophication (Davenport et al., 2005). While fields are typically uniformly fertilized, variable rate precision fertilization offers the potential to reduce expenses while optimizing tuber yield and reducing environmental impact (Ahmad and Sharma, 2023). To enable precision fertilization and to further optimize global resource management, farmers must have data on field yields. Such field yield data can be estimated with above-ground biomass measurements (Liu et al., 2024a, b) or close-range monitoring systems installed on the harvester. Such monitoring systems can automatically estimate key yield parameters such as tuber number, size, volume, and weight. Commonly used monitoring systems use load cells attached to the harvesters’ conveyor belt to measure the mass of the harvested produce in real time (Zamani et al., 2014; Kabir et al., 2018). Such weighing systems are easy to use and maintain, but disadvantages of these systems is that they only capture gross yield and that they include tare (such as stones, soil and crop residue) in the weight measurement. Therefore, there is a need for a tare-free and individualized monitoring of tubers, which allows farmers to learn from their fertilizer management decisions and enable them to collect detailed information on marketable yields for better sales decisions.
@@
Tare-free and individualized monitoring of potato tubers can be accomplished by a camera system and computer vision software. Currently, at least 14 of such camera systems have been presented in scientific literature: Noordam et al. (2000); Hofstee and Molema (2003); ElMasry et al. (2012); Razmjooy et al. (2012); Lee et al. (2018); Long et al. (2018); Si et al. (2018); Su et al. (2018); Pandey et al. (2019); Cai et al. (2020); Lee and Shin (2020); Dolata et al. (2021); Huynh et al. (2022); Jang et al. (2023). Most of these systems (11 of the 14) used a single red, green and blue (RGB) color camera to detect the potato tubers and then estimate their yield parameters using offline calibrated pixel-to-world conversion factors. While this method proved effective in calibrated laboratory setups, it was also found to have limited accuracy for potatoes that are occluded. Occlusion is a common phenomena on the conveyor belt of an operational harvester.
@@
More recent studies have focused on extending the traditional RGB camera with additional three-dimensional (3D) vision. With 3D vision, it is possible to create partial or complete 3D shapes of potato tubers, allowing for better estimation of tuber yield under the challenging conditions of occlusion. Cai et al. (2020) extended an RGB camera with a laser line triangulation method to capture the complete 3D shape of potato tubers. Although their system enabled very accurate volumetric estimates of potato tubers in a laboratory setup, the image acquisition method was unfortunately too time-consuming to be applied on an operational harvester. RGB cameras with embedded depth sensing abilities, known as RGB-D cameras, allow much higher throughput than laser triangulation methods. Although being more susceptible to dust and motion blur, RGB-D cameras are cheaper and easier to integrate into existing machines than laser triangulation methods. RGB-D cameras, such as the ones used by Long et al. (2018) and Su et al. (2018), are therefore more promising for use on an operational harvester.
@@
A current limitation with RGB-D cameras is that they can only produce a partial 3D shape of the potato tuber, which can lead to an underestimation of the actual size, volume or weight. Combining multiple RGB-D cameras can potentially reduce the effect of this problem, but this will not fully solve it and make the overall system complex and more expensive. Therefore, it is more desirable and economically feasible to use a single RGB-D camera, and then estimate the complete 3D shape with dedicated software. There are numerous examples in the scientific literature of using 3D shape completion software to estimate complete shapes from partial shapes. Most of the current shape completion methods use deep learning techniques based on multi-layered perceptrons, graph-based convolutional neural networks and encoder-decoder networks to complete the 3D shape from partially completed point clouds (Fei et al., 2022). Also in the agricultural domain, where deep learning applications are widespread for disease detection, yield prediction and machine automation, there are studies on 3D shape completion: one on the completion of plant leaves (Chen et al., 2023), one on the completion of trees (Xu et al., 2023), and five on the completion of fruits (Ge et al., 2020; Magistri et al., 2022, 2024; Marangoz et al., 2022; Pan et al., 2023).
@@
When assessing the applicability of 3D shape completion methods for potato yield estimation, it is crucial to consider their processing speed. Ideally, the shape completion method should be fast enough to process all the potato tubers that move over the conveyor belt during harvest, because this will provide the most complete and accurate yield estimate possible. Since roughly more than one million potato tubers are harvested per hectare, the 3D shape completion method must complete processing in a matter of milliseconds. Of the studies mentioned above, only the method of Magistri et al. (2022) has the potential to do so. Magistri’s method, called CoRe (Completion and Reconstruction), can complete 3D shapes in 4 milliseconds (tested on sweet peppers and strawberries). The novelty of CoRe is the addition of a convolutional encoder to the deep signed distance function decoder (DeepSDF, Park et al., 2019). CoRe’s convolutional encoder compresses RGB-D images into a latent vector, which is a compact representation of the shape’s geometry. The latent vector from the encoder is then used along with the 3D coordinates of the point cloud as input for DeepSDF to reconstruct the complete 3D shape. Because the latent vector is predicted by the convolutional encoder, there is no need for DeepSDF’s original time-consuming latent vector optimization, making the entire 3D shape completion method significantly faster.
@@
The main contribution of this paper is the extension of the work by Magistri et al. (2022) to answer these research questions. We have conducted a systematic study and analysis on how to optimize the 3D shape completion method. Our optimization has led to several improvements to CoRe’s original convolutional encoder: an improved data preprocessing, geometric and color data augmentations, an updated loss function, an updated neural network architecture, and a graphical processing unit (GPU)-based 3D mesh generation that improves and speeds up the 3D shape completion for precision agriculture applications.
This paper presents the research and development of a 3D shape completion network for estimating the volume of individual potato tubers on an operational harvester. Our research novelties are four-fold. First, we optimized CoRe’s original convolutional encoder for faster and more accurate 3D shape completion of potato tubers on an operational harvester. Second, we performed a systematic analysis of the effect of the latent size, image analysis region, potato size, and potato cultivar on the performance of the 3D shape completion of potato tubers. Third, we conducted two ablation studies on the impact of our new additions to the overall performance. Fourth, we publicly released our 3D dataset with partially and fully completed point clouds of potato tubers collected on an operational harvester in dirty and cluttered circumstances. Although this study focused on potato tubers, our shape completion method can also be applied to other crops that suffer from object occlusion, such as fruits, berries, peppers, tomatoes and other root crops. Our method can thus serve the broader purpose of better handling partial visibility in agricultural applications requiring high throughput, such as robotic harvesting, pick-and-place, path planning, and obstacle detection and avoidance.
@@ 2.1.1 Imaging system
We installed an imaging system above the conveyor belt of a single-row potato harvester (Toyonoki Top-1, Figure 1a). The imaging system was installed in a black plastic box with dimensions of 85 cm (width), 45 cm (depth) and 39 cm (height) (Figure 1b). Inside the plastic box, four light emitting diode (LED) strips were installed to provide light with a color temperature of 6000K (Figure 1c). Our enclosed setup not only ensured uniform lighting conditions, which improved the RGB-D image quality, but also ensured that our system could be used at night. Reflective curtains on the sides of the plastic box helped to diffuse the light. In the center of the box, an RGB-D camera (Intel Realsense D405) was installed that captured images of the conveyor belt with a top-view perspective. The distance between the RGB-D camera and the conveyor belt was approximately 0.33 m. At this distance, the camera’s field of view was 0.64 m (width) by 0.39 m (height). The RGB-D camera was connected to a Lenovo Thinkpad P53 laptop with an Intel Xeon E-2276M 2.8 GHz CPU with 64GB RAM and a Nvidia Quadro RTX 5000 GPU with 16GB memory. On the laptop, the Robot Operating System 2 (ROS2, version: Humble Hawksbill) was used to collect color and depth images with 30 frames per second. The exposure time of the RGB-D camera was set to 5.0 ms, allowing us to capture the images without motion blur. The images were stored in ROS2 bag files.
Figure 1: (a) and (b) give overviews of the imaging system installed on a potato harvester in Sarabetsu, Japan. (c) Inside the imaging box, an RGB-D camera was installed, together with four LED strips that provided the necessary illumination inside the box. The sides of box were covered with a reflective curtain to generate diffuse lighting conditions.
@@ 2.1.2 Image collection on the harvester
While running the harvester in the field or in the barn, we collected images of 339 potato tubers of different sizes and shapes to test our 3D shape completion method. Our collection method consisted of the following procedure: one person standing in front of the imaging box randomly selected a potato tuber and then inserted a colored thumbtack into the potato. The thumbtack was inserted into the potato such that it was visible in the RGB image when the tuber passed the image acquisition region of the RGB-D camera. This procedure was repeated for all 339 potato tubers. Given the speed of the conveyor belt, we were able to capture between 20 and 30 images for each marked potato tuber when it moved under the camera along with the conveyor belt (Figure 2). After the image acquisition, another person standing behind the imaging box collected the marked potato by hand. The marked potato was then placed in a bucket that was later brought to the barn for 3D reconstruction.
@@
Figure 2: Our potato collection method involved marking potato tubers with a colored thumbtack so that the tuber could be easily identified in the image and easily collected after image acquisition. (a), (b), and (c) show a tuber marked with a red thumbtack while it moved over the conveyor belt.
2.1.3 3D reconstruction of potato tubers as ground truth
@@
To obtain the complete 3D shape of the collected potato tubers as ground truth, we have set up a 3D reconstruction system in the barn (Figure 3a). This system consisted of three Canon X7 DSLR cameras, a turntable (Foldio 360, OrangeMonkie Inc.), four auto-detectable marker stands (Fig. 3b), and a photo studio with LED illumination.
@@
Prior to taking the images, each collected potato tuber was pierced with a narrow threaded bolt attached to a white wooden board (Figure 3b). This piercing allowed the tuber to be photographed from three camera perspectives while being held off the ground, allowing almost the entire longitudinal section of the tuber to be photographed at once. To capture all sides of the tuber, the turntable was set to rotate 15 degrees per interval and then stop for two seconds to give the three cameras time to acquire the images. The image acquisition was accomplished with a commercial shutter controller (Esper TriggerBox) that released an electronic trigger to the three cameras at the same time. For each camera, 24 images were recorded per complete rotation of the turntable. The resulting 72 images were directly transferred to the connected laptop computer to store them for 3D reconstruction (Figure 3c). The camera parameter settings, image transferring, and image renaming were accomplished with DigiCamControl software, of which its details are shown in Figure 3d.
Figure 3: The workflow of our 3D reconstruction included three steps: (1) image collection, (2) image preprocessing, (3) 3D reconstruction with Structure-from-Motion (SfM).
After the image acquisition, the images were processed to extract the mask region of the potato tuber. Due to soil residues and light reflections on the tuber surfaces (Figure 3e), which were similar to the black and white background, we had to use a combination of color-based filtering and a deep learning-based segmentation refinement. Given that potato tubers are predominantly yellow (Figure 4), we chose to use the CIELAB color space, because yellow in CIELAB is represented with high positive values on the b* axis and near-zero values on the a* axis, allowing for straightforward filtering without interference from other colors or lighting conditions (we used b* $\geq$ 15 for filtering, see the result in Figure 3f). By using CIELAB filtering as pre-segmentation, it became easier for the deep learning network to refine the mask (Figure 3g). We used the CascadePSP (Cheng et al., 2020) deep learning network due to its multi-stage refinement approach, which enabled better segmentation quality compared to other segmentation networks like U-Net (Ronneberger et al., 2015) and Mask R-CNN (He et al., 2017).
@@
After generating the masks of all tubers, we implemented an automatic batch processing workflow based on the Metashape API for 3D reconstruction (Figure 3i). First, the images and corresponding masks of each potato tuber were grouped for each camera and its corresponding view angles. Then, the control point markers in the images were automatically detected, the scale bars were automatically imported, and world coordinates were automatically assigned to each potato tuber. Afterwards, the images of the different cameras were aligned, and the corresponding key points were generated. These aligned images served as input for the 3D reconstruction, which resulted the colored 3D mesh models for all 339 potato tubers in our dataset. After 3D reconstruction, small distortions, such as 3D meshes belonging to the white wooden board were removed. Then, the filtered meshes were double-checked by another researcher in CloudCompare who removed small disconnected components if necessary. Figure 4 illustrates a filtered 3D mesh of a potato tuber with colored textures.